The length of a rectangle is 3 times its width. If the width is diminished by one meter and its length increased by 3 meters, the area will be increased by 72 sq. meters. What are the original dimensions of the rectangle?
Answer by ankor@dixie-net.com(22740)

The length of a rectangle is 3 times its width.
L = 3W
If the width is diminished by one meter and its length increased by 3 meters, the area will be increased by 72 sq. meters.
(L+3)*(W-1) = LW + 72
Replace L with 3W
(3W+3)*(W-1) = 3W(W) + 72
FOIL the left side
3W^2 - 3W + 3W - 3 = 3W^2 + 72
So you get
3W^2 - 3 = 3W^2 + 72
obviously this will not fly, problem is not solvable

What are the original dimensions of the rectangle?
: